# Parameter Estimation for Weakly Interacting Hypoelliptic Diffusions

### v1

@@ 2.2 Asymptotic Results
    $\displaystyle\bigl{[}\Gamma^{(\mathrm{H})}_{\alpha_{S}}(\theta^{\dagger})\bigr{]}_{k_{1}k_{2}}$ $\displaystyle=4\int_{0}^{T}\mathbb{E}_{\mu_{t}}\Bigl{[}\bigl{(}\partial_{\alpha_{S},k_{1}}V_{S,0}^{\top}\cdot\Sigma_{SS}^{-1}\cdot\partial_{\alpha_{S},k_{2}}V_{S,0}\bigr{)}(\theta^{\dagger},W,\mu_{t})\Bigr{]}dt;$
@@ 4.1 Preliminaries
The first result, Lemma 1, is used and obtained in the works amo:24 ; amo:23 in the sense of point convergence (for fixed $\theta$), where they assumed the global Lipschitz condition on the drift function. We note that Lemma 1 is now established under a locally Lipschitz condition (Assumption 4) on the drift function with the aid of results on finite moments of IPS and McKean-Vlasov SDE, e.g. dos:22 ; dos:19 , which are stated in Lemma 1 in the Supplementary Material supp_AoS . Then, the proof of Lemma 1 is contained in Appendix B.1 in the same material. Lemma 2 also plays a critical role when carrying out asymptotic analysis of the proposed contrast function. The proof of Lemma 2 is provided in Appendix B.2 in the Supplementary Material supp_AoS .
Decomposition of Contrast Function. The fundamental procedure for proving the asymptotic results (Theorems 1, 2) involves decomposing the contrast function (9), scaled by an appropriate factor, into several components. Subsequently, we apply the above two convergence results (Lemmas 1, 2) to these components under Assumptions 1-6. In particular, we often use the following decomposition:
@@ 4.2 Proof of Theorem 1
where $\widetilde{R}^{(4)}\in\mathcal{R}$. We now use Lemmas 1, 2 to obtain that, if $n,N\to\infty$ then ${R}^{(3)}_{n,N}(\theta)\xrightarrow{\mathbb{P}_{\theta^{\dagger}}}\mathcal{Q}(\theta)$, ${R}^{(k)}_{n,N}(\theta)\xrightarrow{\mathbb{P}_{\theta^{\dagger}}}0$, $k=1,2,4,$ uniformly in $\theta\in\Theta$. We thus conclude Lemma 3. ∎
@@
where $f_{\ell_{1}\ell_{2}}^{(1)},\,f_{\ell}^{(2)},f^{(3)}\in\mathcal{R}$ and $h:[0,\infty)\to\mathbb{R}$ such that $|h(t)|\lesssim t$, $t\in[0,\infty)$. Applying Lemmas 1, 2 to $\mathcal{S}_{n,N}^{(\iota),k}(\theta)$, $\iota=1,2,3$, we conclude (24).
@@
where $f_{k_{1}k_{2}}^{(4),\ell},f_{k_{1}k_{2}}^{(5)}\in\mathcal{R}$. We then obtain (27) by applying Lemmas 1, 2 to $\mathcal{T}_{n,N}^{(\iota)}(\theta),\ 1\leq\iota\leq 5$. Due to the uniform convergence w.r.t. $\theta\in\Theta$ and the consistency of $\hat{\alpha}_{S}^{n,N}$, we get (25). Proof of Lemma 4 is now complete. ∎
@@
for $\theta=(\alpha_{S},\alpha_{R},\beta)\in\Theta$, where $f^{(3),k},f^{(4)}\in\mathcal{R}$ and $h^{(\iota)}:[0,\infty)\to\mathbb{R},\,\iota=3,4$, such that $|h^{(\iota)}(t)|\lesssim t$, $t\in[0,\infty)$. We directly have from Lemmas 1, 2 that if $n,N\to\infty$, then: $\mathcal{U}_{n,N}^{(\iota)}(\theta)\xrightarrow{\mathbb{P}_{\theta^{\dagger}}}0,\,\iota=3,4$ and
@@
Assumptions 2-6 together with the finite moments of IPS in $L_{p}$ sense (Lemma 1 in supp_AoS ) leads to that the $\{\widetilde{\mathcal{U}}_{n,N}^{(1)}(\theta)\}_{n,N}$ is a bounded sequence uniformly in $\theta\in\Theta$ in probability. Then, due to (28), we obtain:
@@ 5 Conclusion
This material collects technical results to support the proofs for the main analytics results - Theorems 1 & 2. Appendix A provides some new results to facilitate the analysis in the later sections. Appendix B contains proofs for two key convergence Lemmas 1 & 2 given in Section 4. Appendices C & D show Propositions 1 & 2, respectively, used to prove the CLT in Section 4.3.
@@ A.1 First Result
The results collected in Lemma 1 are also used in the works (amo:24, ; amo:23, ), where Lipschitz conditions are assumed on the drift and diffusion coefficients. In our case, the proof of Lemma 1 is obtained under the weaker locally Lipschitz condition on the drift function, given in Assumption 4.
Proof of Lemma 1. First, (33) follows immediately from (chen:24, , Proposition 2.5) under Assumption 4. We now make use of (33) to get (34) since $\textstyle\mathbb{E}_{\theta}\bigl{[}\mathcal{W}_{p}(\mu_{t}^{N},\delta_{0})^{q}\bigr{]}\leq\tfrac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\theta}|X_{t}^{[i]}|^{q}\leq C$ for some positive constant $C$ independent of $t$ and the index $i$. It remains to show (35). We have that:
@@ A.2 Second Result
We here check (44), and (43) is obtained from a similar argument. First, under Assumptions 1, 3, 4 and 5, we have from Lemma 1 that for any $p\geq 2$, there exists a constant $C>0$ such that for any $\ell_{1},\ell_{2},\ell_{3}=1,\ldots,d,$
@@
It follows from (41), (42) and Lemma 1 that for any $1\leq j\leq n$ and $1\leq i\leq N$, $\mathbb{E}_{\theta^{\dagger}}|D_{j-1}^{[i],k}|^{p}\leq C_{p}\,\Delta_{n}^{p/2}$, $p\geq 1$. We then use Jensen and Cauchy-Schwarz inequalities to get that for any $p\geq 1$, $\mathbb{E}_{\theta^{\dagger}}\bigl{|}\mathbb{E}_{\theta^{\dagger}}\bigl{[}D_{j-1}^{[i_{1}],k_{1}}D_{j-1}^{[i_{2}],k_{2}}|\mathcal{F}_{t_{j-1}}^{N}\bigr{]}\bigr{|}^{p}\leq C_{p}\,\Delta_{n}^{p}.$ Subsequently, we study the rest of terms in the RHS of (51). A further stochastic Taylor expansion for the terms (41) and (42) with Lemma 1 lead to: it holds under Assumptions 1, 3, 4, 5 that $\rho_{S}^{[i]}(j)=\widetilde{\rho}_{S}^{[i]}(j)+\mathcal{E}_{S}^{[i]}(j)$ and $\rho_{R}^{[i]}(j)=\widetilde{\rho}_{R}^{[i]}(j)+\mathcal{E}_{R}^{[i]}(j)$ for some $\mathcal{E}_{S}^{[i]}(j),\mathcal{E}_{R}^{[i]}(j)$ such that $\mathbb{E}_{\theta^{\dagger}}|\mathcal{E}_{S}^{[i]}(j)|^{p}\leq C\Delta_{n}^{5p/2}$ and $\mathbb{E}_{\theta^{\dagger}}|\mathcal{E}_{R}^{[i]}(j)|^{p}\leq C\Delta_{n}^{3p/2}$, $p\geq 1$, where:
@@
We thus conclude Lemma 2.
@@ B.1 Proof of Lemma 1
where we have applied the Cauchy-Schwartz inequality and Lemma 1 – the finite moments of $X^{[i]}_{t}$ uniformly in $t\in[0,T]$ and $\|X_{t_{j-1}}^{[i]}-X_{t}^{[i]}\|_{2}\leq C\Delta_{n}^{1/2}$ for all $t\in[t_{j-1},t_{j}]$ and $1\leq j\leq n$. Thus, $\mathbb{E}_{\theta^{\dagger}}|F_{1}|\to 0$ as $n\to\infty$ and then $F_{1}\xrightarrow{\mathbb{P}_{\theta^{\dagger}}}0$. Similarly, for the second term $F_{2}$, we have that:
@@
where $q(d)$ is a positive number depending on the state dimension $d$ of each particle. In the last inequality, we made use of the following three results: Lemma 1, the time-uniform finite moments bound of McKean-Vlasov SDE under Assumptions 3, 4 (dos:19, , Theorem 3.3), the propagation of chaos (PoC) under Assumptions 3, 4 (dos:22, , Proposition 3.1), specifically, $\sup_{1\leq i\leq N}\mathbb{E}\bigl{[}\sup_{t\in[0,T]}|X_{t}^{[i]}-\bar{X}_{t}^{i}|^{2}\bigr{]}\leq C\times R(N,d),$ with
@@
Also, in the above, to bound the term $\mathcal{W}_{2}(\mu_{t}^{N},\mu_{t})$, we considered: $\|\mathcal{W}_{2}(\mu_{t}^{N},\mu_{t})\|_{2}\leq\|\mathcal{W}_{2}(\mu_{t}^{N},\bar{\mu}^{N}_{t})\|_{2}+\|\mathcal{W}_{2}(\bar{\mu}_{t}^{N},\mu_{t})\|_{2}$, where $\bar{\mu}^{N}_{t}$ is the empirical measure from $N$-independent samples $\{\bar{X}_{t}^{i}\}_{i}$. Then, we have from PoC that $\textstyle\mathbb{E}_{\theta^{\dagger}}[\mathcal{W}_{2}(\mu_{t}^{N},\bar{\mu}_{t}^{N})^{2}]\leq\tfrac{1}{N}\sum_{1\leq i\leq N}\mathbb{E}_{\theta^{\dagger}}|X_{t}^{[i]}-\bar{X}_{t}^{i}|^{2}\leq C\times R(N,d)$ for some constant $C>0$ independent of $t\in[0,T]$. We also obtain that $\mathbb{E}_{\theta^{\dagger}}[\mathcal{W}_{2}(\bar{\mu}_{t}^{N},\mu_{t})^{2}]\leq C\times R(N,d)$ by applying the result (carmona:18, , Theorem 5.8) with the time-uniform finite moments of McKean-Vlasov SDE. Thus, we have $\mathbb{E}_{\theta^{\dagger}}|F_{2}|\xrightarrow{\mathbb{P}_{\theta^{\dagger}}}0$, and we conclude the pointwise convergence of (17). To prove the uniform convergence w.r.t $\theta\in\Theta$, it sufficies to show that the sequence of $C(\Theta;\mathbb{R})$-valued random variable $\textstyle\bigl{\{}\mathcal{S}_{n,N}(\theta)\bigr{\}}_{n,N}\equiv\bigl{\{}\tfrac{\Delta_{n}}{N}\sum_{i=1}^{N}\sum_{j=1}^{n}f(\theta,X_{t_{j-1}}^{N},\mu_{t_{j-1}}^{N})\bigr{\}}_{n,N}$ is tight. Under the property (16) of the function $f\in\mathcal{R}$ together with Lemma 1, it holds that for all $n,N$:

### v2

@@ 2.2 Asymptotic Results
Assumption 7 is related to the identifiability of the statistical parameters. A similar condition is assumed in DeH23 where the authors studied the local asymptotic normality (LAN) for an $N$ weakly interacting elliptic diffusions under the continuous observation regime and discussed identifiability and non-degeneracy of asymptotic Fisher information. Assumption 7 is used in proving the consistency of the contrast estimator in Section 4.2 and also implies the uniqueness of the minimiser for the objective function (9) for large enough $n,N$; see Remark 8.
@@
    $\displaystyle\bigl[\Gamma^{(\mathrm{H})}_{\alpha_{S}}(\theta^{\dagger})\bigr]_{k_{1}k_{2}}$ $\displaystyle=4\int_{0}^{T}\mathbb{E}_{\mu_{t}}\Bigl[\bigl(\partial_{\alpha_{S},k_{1}}V_{S,0}^{\top}\cdot\Sigma_{SS}^{-1}\cdot\partial_{\alpha_{S},k_{2}}V_{S,0}\bigr)(\theta^{\dagger},W,\mu_{t})\Bigr]dt;$ (10)
@@
Remark 5.
[…]
$\displaystyle V_{S,0}(\alpha_{S},x)=x_{R},\quad V_{R,j}(\beta,x,\mu)=\beta_{j}e_{j},\quad j=1,\ldots,d_{B},$
with $\beta=(\beta_{1},\ldots,\beta_{d_{B}})\in\Theta_{\beta},\,d_{\beta}=d_{B}$, for $x=(x_{S},x_{R})\in\mathbb{R}^{d}$, $\alpha_{S}\in\Theta_{\alpha_{S}}$, $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, where $e_{j}$ is the standard unit vector. Then, we have:
$\displaystyle\Gamma^{(\mathrm{H})}_{\beta}(\theta^{\dagger})=4T\times\mathrm{diag}\bigl[(\beta_{1}^{\dagger})^{-2},\ldots,(\beta_{d_{B}}^{\dagger})^{-2}\bigr]\in\mathbb{R}^{d_{B}\times d_{B}},$
which is invertible as long as $\beta^{\dagger}_{j}\neq 0$, $j=1,\ldots,d_{B}$. For the precision matrices (10), (11), associated with the drift parameter $\alpha=(\alpha_{S},\alpha_{R})$, one may be able to derive a sufficient condition for these to be positive definite under Assumption 2, similarly to Proposition 15 in DeH23 . Here, let us focus on (10) and assume for simplicity that $d_{\alpha_{S}}=1$. Then, noticing that $\Sigma_{SS}^{-1}$ is positive definite, we have that $\partial_{\alpha_{S}}V_{S,0}^{\top}(\alpha_{S}^{\dagger},x)\cdot\,\Sigma_{SS}^{-1}(\theta^{\dagger},x,\mu)\,\cdot\partial_{\alpha_{S}}V_{S,0}(\alpha_{S}^{\dagger},x)\geq 0$ for all $(x,\theta)\in\mathbb{R}^{d}\times\mathcal{P}_{2}(\mathbb{R}^{d})$ where $0$ is attained iff $\partial_{\alpha_{S}}V_{S,0}(\alpha_{S}^{\dagger},x)=\mathbf{0}_{d_{S}}$. This implies that $\Gamma_{\alpha_{R}}^{(\mathrm{H})}(\theta^{\dagger})>0$ if there exists $(A,B)\in\mathcal{B}([0,T])\otimes\mathcal{B}(\mathbb{R}^{d})$ with $\mathrm{Leb}(A)>0$ s.t. for all $t\in A$, we have $\mu_{t}(B)>0$ and $\partial_{\alpha_{S}}V_{S,0}(\alpha_{S}^{\dagger},x)\neq\mathbf{0}_{d_{S}}$, $x\in B$. Similar arguments can be applied for the case of $\alpha_{R}$. Further exploration under more general settings is left for future research.
@@ 4.1 Preliminaries
The first result, Lemma 2, is used and obtained in the works amo:24 ; amo:23 in the sense of point convergence (for fixed $\theta$), where they assumed the global Lipschitz condition on the drift function. We note that Lemma 2 is now established under a locally Lipschitz condition (Assumption 4) on the drift function with the aid of results on finite moments of IPS and McKean-Vlasov SDE, e.g. dos:22 ; dos:19 , which are stated in Lemma A.1 in the Supplementary Material supp_AoS . Then, the proof of Lemma 2 is contained in Appendix B.1 in the same material. Lemma 3 also plays a critical role when carrying out asymptotic analysis of the proposed contrast function. The proof of Lemma 3 is provided in Appendix B.2 in the Supplementary Material supp_AoS .
Decomposition of Contrast Function. The fundamental procedure for proving the asymptotic results (Theorems 1, 2) involves decomposing the contrast function (9), scaled by an appropriate factor, into several components. Subsequently, we apply the above two convergence results (Lemmas 2, 3) to these components under Assumptions 1-6. In particular, we often use the following decomposition:
@@ 4.2 Proof of Theorem 1
where $\widetilde{R}^{(4)}\in\mathcal{R}$. We now use Lemmas 2, 3 to obtain that, if $n,N\to\infty$ then ${R}^{(3)}_{n,N}(\theta)\xrightarrow{\mathbb{P}_{\theta^{\dagger}}}\mathcal{Q}(\theta)$, ${R}^{(k)}_{n,N}(\theta)\xrightarrow{\mathbb{P}_{\theta^{\dagger}}}0$, $k=1,2,4,$ uniformly in $\theta\in\Theta$. We thus conclude Lemma 4. ∎
@@
Remark 8.
Lemma 4 implies that optimising the scaled objective function $\tfrac{\Delta_{n}}{N}\ell_{n,N}(\theta)$ is asymptotically equivalent to optimising function $\mathcal{Q}(\theta)$ given in (23). In particular, $Q(\theta)$ admits the unique minima $\alpha_{S}=\alpha_{S}^{\dagger}$ under Assumptions 2 and 7 due to the positive definiteness of $\Lambda_{SS}(\theta,W,\mu^{\dagger})$ for any $(\theta,W)\in\Theta\times\mathbb{R}^{d}$ and $b_{S}(\alpha_{S},x)\neq 0$, $\mu_{t}(dx)\,\mathrm{a.e.}\,$ for all $t\in[0,T]$ whenever $\alpha_{S}\neq\alpha_{S}^{\dagger}$. Similarly, the asymptotic results below imply that scaled objective functions asymptotically allow a unique minimum for the cases of $\alpha_{R}$ and $\beta$ as well under Assumptions 2 and 7.
@@
where $f_{\ell_{1}\ell_{2}}^{(1)},\,f_{\ell}^{(2)},f^{(3)}\in\mathcal{R}$ and $h:[0,\infty)\to\mathbb{R}$ such that $|h(t)|\lesssim t$, $t\in[0,\infty)$. Applying Lemmas 2, 3 to $\mathcal{S}_{n,N}^{(\iota),k}(\theta)$, $\iota=1,2,3$, we conclude (25).
@@
where $f_{k_{1}k_{2}}^{(4),\ell},f_{k_{1}k_{2}}^{(5)}\in\mathcal{R}$. We then obtain (28) by applying Lemmas 2, 3 to $\mathcal{T}_{n,N}^{(\iota)}(\theta),\ 1\leq\iota\leq 5$. Due to the uniform convergence w.r.t. $\theta\in\Theta$ and the consistency of $\hat{\alpha}_{S}^{n,N}$, we get (26). Proof of Lemma 5 is now complete. ∎
@@
for $\theta=(\alpha_{S},\alpha_{R},\beta)\in\Theta$, where $f^{(3),k},f^{(4)}\in\mathcal{R}$ and $h^{(\iota)}:[0,\infty)\to\mathbb{R},\,\iota=3,4$, such that $|h^{(\iota)}(t)|\lesssim t$, $t\in[0,\infty)$. We directly have from Lemmas 2, 3 that if $n,N\to\infty$, then: $\mathcal{U}_{n,N}^{(\iota)}(\theta)\xrightarrow{\mathbb{P}_{\theta^{\dagger}}}0,\,\iota=3,4$ and
@@
Assumptions 2-6 together with the finite moments of IPS in $L_{p}$ sense (Lemma A.1 in supp_AoS ) leads to that the $\{\widetilde{\mathcal{U}}_{n,N}^{(1)}(\theta)\}_{n,N}$ is a bounded sequence uniformly in $\theta\in\Theta$ in probability. Then, due to (29), we obtain:
@@ 5 Conclusion
This material collects technical results to support the proofs for the main analytics results - Theorems 1 & 2. Appendix A provides some new results to facilitate the analysis in the later sections. Appendix B contains proofs for two key convergence Lemmas 2 & 3 given in Section 4. Appendices C & D show Propositions 1 & 2, respectively, used to prove the CLT in Section 4.3.
@@ A.1 First Result
The results collected in Lemma A.1 are also used in the works (amo:24, ; amo:23, ), where Lipschitz conditions are assumed on the drift and diffusion coefficients. In our case, the proof of Lemma A.1 is obtained under the weaker locally Lipschitz condition on the drift function, given in Assumption 4.
Proof of Lemma A.1. First, (34) follows immediately from (chen:24, , Proposition 2.5) under Assumption 4. We now make use of (34) to get (35) since $\textstyle\mathbb{E}_{\theta}\bigl[\mathcal{W}_{p}(\mu_{t}^{N},\delta_{0})^{q}\bigr]\leq\tfrac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\theta}|X_{t}^{[i]}|^{q}\leq C$ for some positive constant $C$ independent of $t$ and the index $i$. It remains to show (36). We have that:
@@ A.2 Second Result
We here check (45), and (44) is obtained from a similar argument. First, under Assumptions 1, 3, 4 and 5, we have from Lemma A.1 that for any $p\geq 2$, there exists a constant $C>0$ such that for any $\ell_{1},\ell_{2},\ell_{3}=1,\ldots,d,$
@@
It follows from (42), (43) and Lemma A.1 that for any $1\leq j\leq n$ and $1\leq i\leq N$, $\mathbb{E}_{\theta^{\dagger}}|D_{j-1}^{[i],k}|^{p}\leq C_{p}\,\Delta_{n}^{p/2}$, $p\geq 1$. We then use Jensen and Cauchy-Schwarz inequalities to get that for any $p\geq 1$, $\mathbb{E}_{\theta^{\dagger}}\bigl|\mathbb{E}_{\theta^{\dagger}}\bigl[D_{j-1}^{[i_{1}],k_{1}}D_{j-1}^{[i_{2}],k_{2}}|\mathcal{F}_{t_{j-1}}^{N}\bigr]\bigr|^{p}\leq C_{p}\,\Delta_{n}^{p}.$ Subsequently, we study the rest of terms in the RHS of (52). A further stochastic Taylor expansion for the terms (42) and (43) with Lemma A.1 lead to: it holds under Assumptions 1, 3, 4, 5 that $\rho_{S}^{[i]}(j)=\widetilde{\rho}_{S}^{[i]}(j)+\mathcal{E}_{S}^{[i]}(j)$ and $\rho_{R}^{[i]}(j)=\widetilde{\rho}_{R}^{[i]}(j)+\mathcal{E}_{R}^{[i]}(j)$ for some $\mathcal{E}_{S}^{[i]}(j),\mathcal{E}_{R}^{[i]}(j)$ such that $\mathbb{E}_{\theta^{\dagger}}|\mathcal{E}_{S}^{[i]}(j)|^{p}\leq C\Delta_{n}^{5p/2}$ and $\mathbb{E}_{\theta^{\dagger}}|\mathcal{E}_{R}^{[i]}(j)|^{p}\leq C\Delta_{n}^{3p/2}$, $p\geq 1$, where:
@@
We thus conclude Lemma A.2.
@@ B.1 Proof of Lemma 2
where we have applied the Cauchy-Schwartz inequality and Lemma A.1 – the finite moments of $X^{[i]}_{t}$ uniformly in $t\in[0,T]$ and $\|X_{t_{j-1}}^{[i]}-X_{t}^{[i]}\|_{2}\leq C\Delta_{n}^{1/2}$ for all $t\in[t_{j-1},t_{j}]$ and $1\leq j\leq n$. Thus, $\mathbb{E}_{\theta^{\dagger}}|F_{1}|\to 0$ as $n\to\infty$ and then $F_{1}\xrightarrow{\mathbb{P}_{\theta^{\dagger}}}0$. Similarly, for the second term $F_{2}$, we have that:
@@
where $q(d)$ is a positive number depending on the state dimension $d$ of each particle. In the last inequality, we made use of the following three results: Lemma A.1, the time-uniform finite moments bound of McKean-Vlasov SDE under Assumptions 3, 4 (dos:19, , Theorem 3.3), the propagation of chaos (PoC) under Assumptions 3, 4 (dos:22, , Proposition 3.1), specifically, $\sup_{1\leq i\leq N}\mathbb{E}\bigl[\sup_{t\in[0,T]}|X_{t}^{[i]}-\bar{X}_{t}^{i}|^{2}\bigr]\leq C\times R(N,d),$ with
@@
Also, in the above, to bound the term $\mathcal{W}_{2}(\mu_{t}^{N},\mu_{t})$, we considered: $\|\mathcal{W}_{2}(\mu_{t}^{N},\mu_{t})\|_{2}\leq\|\mathcal{W}_{2}(\mu_{t}^{N},\bar{\mu}^{N}_{t})\|_{2}+\|\mathcal{W}_{2}(\bar{\mu}_{t}^{N},\mu_{t})\|_{2}$, where $\bar{\mu}^{N}_{t}$ is the empirical measure from $N$-independent samples $\{\bar{X}_{t}^{i}\}_{i}$. Then, we have from PoC that $\textstyle\mathbb{E}_{\theta^{\dagger}}[\mathcal{W}_{2}(\mu_{t}^{N},\bar{\mu}_{t}^{N})^{2}]\leq\tfrac{1}{N}\sum_{1\leq i\leq N}\mathbb{E}_{\theta^{\dagger}}|X_{t}^{[i]}-\bar{X}_{t}^{i}|^{2}\leq C\times R(N,d)$ for some constant $C>0$ independent of $t\in[0,T]$. We also obtain that $\mathbb{E}_{\theta^{\dagger}}[\mathcal{W}_{2}(\bar{\mu}_{t}^{N},\mu_{t})^{2}]\leq C\times R(N,d)$ by applying the result (carmona:18, , Theorem 5.8) with the time-uniform finite moments of McKean-Vlasov SDE. Thus, we have $\mathbb{E}_{\theta^{\dagger}}|F_{2}|\xrightarrow{\mathbb{P}_{\theta^{\dagger}}}0$, and we conclude the pointwise convergence of (18). To prove the uniform convergence w.r.t $\theta\in\Theta$, it sufficies to show that the sequence of $C(\Theta;\mathbb{R})$-valued random variable $\textstyle\bigl\{\mathcal{S}_{n,N}(\theta)\bigr\}_{n,N}\equiv\bigl\{\tfrac{\Delta_{n}}{N}\sum_{i=1}^{N}\sum_{j=1}^{n}f(\theta,X_{t_{j-1}}^{N},\mu_{t_{j-1}}^{N})\bigr\}_{n,N}$ is tight. Under the property (17) of the function $f\in\mathcal{R}$ together with Lemma A.1, it holds that for all $n,N$: